Probability and Statistics II GSIA, Carnegie Mellon University

45-734, Spring 2002

Homework #1, due Thursday, March 21, 2002

Please do the following problems from the book:

pg 439  3, 5
pp 455-8 13, 17, 25 (don’t forget answer to 16 in back of book)